OpenStudy (anonymous):

A cannery processed 1,445 pounds of strawberries in 4.5 hours. The cannery processed 2,320 pounds in 7 hours. Write a linear equation to model the relationship between the weight of strawberries S and time T. How many pounds of strawberries can be processed in 11 hours?

OpenStudy (anonymous):

First we need to find the slope so we need to write this info as 2 points. Which variable, s or t, would be the dependent variable?

OpenStudy (anonymous):

T?

OpenStudy (anonymous):

Nope, it would be s. In other words the weight of the strawberries produced depends on the time

OpenStudy (anonymous):

Actually S since it depends on the weight...right?

OpenStudy (anonymous):

S depends on the value of t. so s would be your 'y' and t would be your 'x' in (x,y)

OpenStudy (anonymous):

@CalebBeavers could you please help me on my question!?

OpenStudy (anonymous):

Okay, so what do we do now that we know that?

OpenStudy (anonymous):

Make two points with the info given. They would be (9/2, 1445) and (7, 2320)

OpenStudy (anonymous):

Now find the slope of those two points

OpenStudy (anonymous):

-2.5/875

OpenStudy (anonymous):

*2.5

OpenStudy (anonymous):

So it that the amount for one hour?

OpenStudy (anonymous):

*is

OpenStudy (anonymous):

you got the top and bottom mixed up, it should be 875/2.5 which is 350

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

so now that we have a slope we can put this into point slope form and simplify it which will give us y=mx+b

OpenStudy (anonymous):

Have you ever done point slope form?

OpenStudy (anonymous):

Yea, but I basically suck at it

OpenStudy (anonymous):

Ha thats alright, ill walk you through it. the formula is \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

Where m is the slope, and (x1, y1) is whichever point you wanna use

OpenStudy (anonymous):

So which one?

OpenStudy (anonymous):

7 I guess lol

OpenStudy (anonymous):

Alright, (7, 2320). when we plug that and the slope in we get: y-2320=350(x-7)

OpenStudy (anonymous):

Distribute the 350 and what do you get?

OpenStudy (anonymous):

We're dividing 350 by both sides?

OpenStudy (anonymous):

no distribute it into the parentheses

OpenStudy (anonymous):

<~< No clue how to do that (just started learning this today so bare with me)

OpenStudy (anonymous):

Where you multiply 350 times everything in the parentheses, so you'd get: y-2320=350x-2450

OpenStudy (anonymous):

ohhh I feel dumb lol

OpenStudy (anonymous):

haha its alright, now that weve done that, add the 2320 to both sides

OpenStudy (anonymous):

2320y=350x-130. right? (otherwise I'm guessing 2320y=2670x-130)

OpenStudy (anonymous):

You should only be left with y on the left, when you add 2320 it gets rid of it on the left and then you just add it to the last number on the right

OpenStudy (anonymous):

x.x oops

OpenStudy (anonymous):

Soo, it's actually y=350x-130?

OpenStudy (anonymous):

yep. And since we said s is the y variable and t is the x variable, lets go ahead and change those. so we have: s=350t-130 so if you wanna find how many pounds you can produce in 11 hours, what would you do?

OpenStudy (anonymous):

You would substitute and get s=350(11)-130 or y=3720

OpenStudy (anonymous):

Yep so you can produce 3720 pounds in 11 hours

OpenStudy (anonymous):

Wow that's super easy actually , thanks so much!

OpenStudy (anonymous):

You're welcome